OpenStudy (genny7):

Explain how to solve one absolute value equation by writing it as two equations

OpenStudy (cwrw238):

an example is solve for x |x - 2| = 4 this means that x - 2 = 4 or x - 2 = -4 solve these 2 for x

OpenStudy (cwrw238):

ok?

OpenStudy (genny7):

yes I got thank you very much

OpenStudy (cwrw238):

yw